OpenStudy (anonymous):

Frank has the face cards and aces of a deck of cards. How many permutations and combinations are there of these cards taken 4 at a time? (Remember that a deck of cards has 4 suits each containing an ace, king, queen, and jack.)

OpenStudy (anonymous):

@jim_thompson5910 I know I use these.

OpenStudy (kropot72):

There are 12 face cards and 4 aces giving 16 different cards. Just substitute the values n = 16 and r = 4 in the formulas. \[16P4=\frac{16!}{(16-4)!}=?\] \[16C4=\frac{16!}{4!(16-4)!}=?\]

OpenStudy (anonymous):

I don't know how to solve those equations...

OpenStudy (kropot72):

Do you know what a factorial is? For example 3! = 3 * 2 * 1 and 5! = 5 * 4 * 3 * 2 * 1. So 3! = 6 and 5! = 120 You might have a calculator that has the factorial function. Look for the exclamation sign.

OpenStudy (anonymous):

Well 16! come up as 2.09228989E13

OpenStudy (anonymous):

And 4!=24

OpenStudy (kropot72):

Good work! My calculator gives the same results. Keep in mind that (16 - 4)! = 12! So for the permutations find 16! and divide by 12!.

OpenStudy (anonymous):

So 43680

OpenStudy (kropot72):

Correct! There are 43680 permutations of the 16 cards taken 4 at a time. Now divide that answer by 4! to find the combinations>

OpenStudy (anonymous):

10920

OpenStudy (anonymous):

It was wrong.

OpenStudy (anonymous):

Divide by 4!

OpenStudy (kropot72):

Yes divide the permutations result by 4! to find the combinations.

OpenStudy (anonymous):

You said 4. So I got the answer wrong the first time. Thanks for the help!

OpenStudy (kropot72):

Your welcome :) BTW the first posting for combinations calculation shows ".....divide that answer by 4! ......." on my computer.